# Client-Centered Federated Learning for Heterogeneous EHRs: Use Fewer Participants to Achieve the Same Performance

### v4

@@ 4.2 Relationship between Averaged Patient Embedding Similarity and Host Performance
Table 3: Correlation between host-subject similarity and the change in host performance when a subject participates in federated learning compared to when the subject does not. The correlations are averaged across that of FedAvg, FedProx, FedBN, and FedPxN. Similarity is measured using averaged patient embeddings (or averaged softmax for KL divergence). See Appendix D for details.
@@
Table 3 presents the observed correlations between (1) similarity between the host and a subject, measured using averaged patient embeddings (or averaged softmax values for KL Divergence) and (2) the difference in host performance with and without the subject participating in federated learning (if multiple subjects participate, other subjects consistently participate in both scenarios to isolate the effect of the target subject). The results show consistent correlations across all three similarity methods and the different numbers of participating clients. This trend holds across each of the four federated learning algorithms (see Appendix D). Note that Cosine Similarity exhibits positive correlations (as higher values indicate greater similarity), while Euclidean Distance and KL Divergence exhibit negative correlations (as higher values indicate lower similarity). These correlations indicate that averaged patient embedding similarity can serve as a reliable criterion for selecting which subjects to exclude from federated learning, particularly by eliminating those with lower similarity scores (or larger distances).
@@ 4.3 Selecting Participating Subjects using Averaged Patient Embedding Similarity
Table 4: Host performance (average macro AUROC across 12 tasks) when selecting participating subjects using averaged patient embeddings (Cosine Similarity, Euclidean Distance, and KL Divergence), averaged across FedAvg, FedProx, FedBN, FedPxN. The performance is compared to (1) using all 5 clients, (2) Average, and (3) Best. Average and Best refer to the average and best host performance, respectively, across all subject combinations for a fixed number of participating clients. Underlined values indicate performance equal to or better than when using all 5 clients, bold indicates performance equal to Best performance, asterisk indicates performance exceeding Average performance. Standard deviations are listed in Appendix E.
@@
Detailed performance results for each individual federated learning algorithm are presented in Appendix F (see Tables 11 and 12). These results show that our selection method consistently achieves performance equal to or better than using all 5 clients and consistently matches the Best and exceeds the Average performance in most cases across each of the four algorithms. Overall, these results highlight the effectiveness of our averaged patient embedding-based method in identifying suitable subjects for the host, thereby enabling effective reduction in the number of participating clients without sacrificing host model performance.
@@ Discussion
In this paper, we introduce EHRFL, a federated learning framework for developing a client-specific model using EHRs. By incorporating text-based EHR modeling, EHRFL enables federated learning across clients with heterogeneous EHR systems. Furthermore, our framework optimizes the selection of participating clients using averaged patient embeddings, reducing the number of participants without compromising model performance. Our experiments demonstrate the efficacy of EHRFL in enabling federated learning across distinct EHR systems, while also reducing costs by identifying suitable participants. One limitation of this study is that our experiments were conducted solely on ICU datasets. While is primarily due to the limited availability of open-source EHR data, in future works, we plan to evaluate our framework on more diverse EHR systems to further validate its generalizability. Overall, we believe our framework provides a practical solution for leveraging federated learning in healthcare, providing a pathway for cost-effective client-specific model development.
Acknowledgements
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant (No.RS-2019-II190075), and National Research Foundation of Korea (NRF) grant (NRF-2020H1D3A2A03100945), funded by the Korea government (MSIT).
Author contributions statement
J.K.: Designed the method, conducted experiments, and wrote the manuscript. J.K.: provided discussions and feedback, and reviewed the manuscript. K.H.: provided discussions. E.C.: provided supervisions and analysis, and reviewed the manuscript.
Data availability
This paper uses MIMIC-III [21], MIMIC-IV [22], and eICU [23] datasets, which are available on Physionet [26] with Physionet credentials.
Code availability
The source code used in this research is publicly available on Github: https://github.com/ji-youn-kim/EHRFL.
Additional information
The author(s) declare no competing interests.
Appendix A Prediction Tasks
Table 5: Prediction Tasks and Descriptions
Task Name Task Type Description
Mortality (Mort) Binary Predict whether the patient will die within the next 24 hours.
Two-Week Mortality (Mort_2w) Binary Predict whether the patient will die within the next two weeks.
Three-Day ICU Stay (Los_3day) Binary Predict whether the patient’s ICU stay will exceed three days.
Seven-Day ICU Stay (Los_7day) Binary Predict whether the patient’s ICU stay will exceed seven days.
ICU Readmission (Readm) Binary Predict whether the patient will be readmitted to the ICU during the same hospital admission.
Bilirubin Level (Bilirubin) Multi-Class Predict serum bilirubin levels, categorized into five discrete classes.
Creatinine Level (Creatinine) Multi-Class Predict serum creatinine levels, categorized into five discrete classes.
Discharge Acuity (Final Acuity) Multi-Class Predict the patient’s discharge location at the end of the hospital stay.
Imminent Discharge (Im_Disch) Multi-Class Predict whether the patient will be discharged within the next 24 hours, along with the discharge location.
Platelet Count (Platelets) Multi-Class Predict platelet counts, categorized into five discrete classes.
White Blood Cell Count (WBC) Multi-Class Predict white blood cell counts, categorized into three discrete classes.
Diagnosis Prediction (Diagnosis) Multi-Label Predict all diagnosis codes associated with the patient’s admission, classified into 18 categories.
Appendix B Client Statistics
Table 6: Demographic Information of the Five Clients.
MIMIC-III-MV MIMICIII-CV MIMIC-IV eICU-West eICU-South
Cohort Size 20,288 24,710 34,888 27,656 45,379
Gender (%) Male 55.8 56.5 55.8 54.9 53.1
Female 44.2 43.5 44.2 45.1 46.9
Ethnicity (%) White 72.8 70.6 66.0 77.2 68.5
Black 10.6 8.7 10.9 4.6 21.0
Hispanic 4.1 2.9 3.6 4.9 5.4
Asian 2.7 2.0 3.1 3.2 1.1
Others 9.8 15.8 16.4 10.1 4.0
Table 7: Label distributions of the Five Clients.
Type Task Label MIMIC-III-MV (%) MIMIC-III-CV (%) MIMIC-IV (%) eICU-West (%) eICU-South (%)
Binary Mort True 1.6 1.8 1.6 1.8 1.8
Mort_2w True 7.7 9.0 7.8 7.3 7.9
Los_3day True 30.9 36.7 29.9 25.5 29.9
Los_7day True 10.5 13.4 9.0 7.9 9.1
Readm True 5.6 5.4 7.3 14.6 8.7
Multi-Class Bilirubin 0 14.7 12.0 18.1 26.8 17.5
1 3.0 2.7 3.7 4.2 2.7
2 4.2 3.1 4.1 3.3 2.1
3 1.2 0.9 1.1 0.6 0.4
4 0.8 0.8 0.9 0.5 0.2
Creatinine 0 58.1 60.0 58.9 50.8 47.6
1 16.2 16.0 16.5 13.4 15.6
2 6.1 6.2 6.5 5.9 7.2
3 1.5 1.6 1.8 1.8 2.5
4 1.3 1.4 1.8 2.0 2.7
Final Acuity 0 51.5 52.3 52.5 61.1 61.2
1 3.1 3.6 3.8 3.1 3.9
2 6.4 7.8 6.1 5.3 5.3
3 13.2 6.7 10.6 12.1 15.4
4 8.4 15.7 6.7 1.9 4.6
5 17.5 13.9 19.5 15.3 9.1
Im_Disch 0 1.6 1.8 1.6 1.8 1.8
1 2.6 1.9 2.8 5.6 4.9
2 95.0 95.7 94.4 90.7 91.8
3 0.7 0.4 0.5 1.7 1.3
4 0.0 0.1 0.0 0.0 0.0
5 0.1 0.1 0.1 0.1 0.1
Platelets 0 61.7 61.9 55.6 45.6 48.9
1 16.4 17.3 21.0 15.1 15.9
2 7.2 7.7 9.4 7.2 7.2
3 2.0 1.4 2.5 1.8 1.6
4 0.3 0.2 0.4 0.3 0.3
WBC 0 4.3 3.1 3.5 2.4 2.6
1 55.9 51.1 54.1 43.8 44.4
2 27.2 33.8 31.3 26.9 26.9
Multi-Label Diagnosis 0 33.6 26.1 38.4 18.1 10.4
1 31.1 22.6 32.1 5.4 6.0
2 78.0 65.3 83.1 24.1 31.4
3 46.3 33.0 57.6 7.8 13.2
4 48.0 28.2 68.8 10.8 11.0
5 41.6 23.0 52.0 14.6 15.6
6 85.8 82.5 89.2 51.1 57.6
7 52.7 48.2 54.7 32.9 34.2
8 53.3 39.8 58.6 14.0 14.6
9 52.1 39.9 55.6 18.1 22.4
10 0.4 0.4 0.6 0.4 0.6
11 12.0 9.8 11.8 1.8 1.4
12 31.5 16.0 35.3 1.6 1.6
13 3.9 3.0 5.1 0.0 0.1
14 46.3 43.7 43.2 13.7 13.8
15 39.4 20.2 59.5 11.7 13.0
16 59.4 33.1 75.7 3.9 7.2
Appendix C Standard Deviation Results for Text-based EHR Modeling
Table 8: Standard Deviations for Table 2. The values represent standard deviations computed across three random seeds, each with distinct data splits and model initializations.
1 Single 0.004 0.004 0.007 0.004 0.005
- - Average Best Average Best Average Best Average Best Average Best
2 FedAvg 0.003 0.004 0.005 0.004 0.004 0.003 0.005 0.006 0.003 0.003
FedProx 0.004 0.004 0.003 0.003 0.005 0.005 0.006 0.007 0.003 0.002
FedBN 0.002 0.002 0.005 0.006 0.004 0.004 0.006 0.007 0.005 0.005
FedPxN 0.005 0.005 0.003 0.005 0.005 0.005 0.005 0.008 0.004 0.005
Avg. 0.004 0.004 0.004 0.005 0.005 0.004 0.006 0.007 0.004 0.004
3 FedAvg 0.005 0.003 0.004 0.004 0.002 0.003 0.006 0.005 0.005 0.005
FedProx 0.004 0.002 0.004 0.005 0.002 0.001 0.005 0.005 0.005 0.006
FedBN 0.004 0.002 0.004 0.002 0.003 0.004 0.004 0.003 0.006 0.006
FedPxN 0.002 0.002 0.004 0.003 0.003 0.003 0.004 0.006 0.004 0.007
Avg. 0.004 0.002 0.004 0.004 0.003 0.003 0.005 0.005 0.005 0.006
4 FedAvg 0.004 0.004 0.004 0.003 0.003 0.004 0.008 0.007 0.008 0.008
FedProx 0.003 0.001 0.006 0.003 0.001 0.002 0.005 0.005 0.006 0.006
FedBN 0.005 0.006 0.005 0.005 0.002 0.002 0.007 0.006 0.006 0.007
FedPxN 0.003 0.003 0.004 0.004 0.001 0.004 0.005 0.006 0.005 0.006
Avg. 0.004 0.004 0.005 0.004 0.002 0.003 0.006 0.006 0.006 0.007
5 FedAvg 0.006 0.006 0.007 0.007 0.002 0.002 0.008 0.008 0.008 0.008
FedProx 0.006 0.006 0.006 0.006 0.003 0.003 0.005 0.005 0.004 0.004
FedBN 0.007 0.007 0.004 0.004 0.002 0.002 0.006 0.006 0.008 0.008
FedPxN 0.006 0.006 0.003 0.003 0.003 0.003 0.004 0.004 0.009 0.009
Avg. 0.006 0.006 0.005 0.005 0.003 0.003 0.006 0.006 0.007 0.007
Appendix D Correlation Values for each Federated Learning Algorithm
Table 9: Correlation between host-subject similarity and host performance change with and without subject participation in FL. Similarity is measured using averaged patient embeddings. For each of the 4 algorithms, the number of data points are 60 for 2 and 5 participating clients, and 180 for 3 and 4 participating clients. Breakdown: 2 clients: 5 hosts × 4 target subjects × 3 seeds; 3 clients: 5 hosts × 4 target subjects × 3 combinations (1 additional client) × 3 seeds; 4 clients: 5 hosts × 4 target subjects × 3 combinations (2 additional clients) × 3 seeds; 5 clients: 5 hosts × 4 target subjects × 1 (3 additional clients) x 3 seeds.
FedAvg
Method Metric 2 Clients 3 Clients 4 Clients 5 Clients Overall
Cosine Spearman 0.381 0.411 0.389 0.494 0.402
Kendall 0.266 0.278 0.271 0.375 0.277
Euclidean Spearman -0.416 -0.425 -0.385 -0.460 -0.407
Kendall -0.289 -0.289 -0.269 -0.340 -0.279
KL Spearman -0.455 -0.391 -0.391 -0.377 -0.393
Kendall -0.321 -0.266 -0.268 -0.271 -0.269
FedProx
Method Metric 2 Clients 3 Clients 4 Clients 5 Clients Overall
Cosine Spearman 0.526 0.546 0.476 0.385 0.489
Kendall 0.376 0.388 0.325 0.269 0.341
Euclidean Spearman -0.574 -0.535 -0.485 -0.356 -0.494
Kendall -0.401 -0.377 -0.331 -0.253 -0.342
KL Spearman -0.623 -0.506 -0.480 -0.337 -0.484
Kendall -0.438 -0.360 -0.327 -0.237 -0.339
FedBN
Method Metric 2 Clients 3 Clients 4 Clients 5 Clients Overall
Cosine Spearman 0.431 0.383 0.350 0.434 0.382
Kendall 0.310 0.266 0.247 0.302 0.267
Euclidean Spearman -0.457 -0.395 -0.313 -0.397 -0.372
Kendall -0.322 -0.271 -0.218 -0.270 -0.256
KL Spearman -0.469 -0.397 -0.295 -0.380 -0.366
Kendall -0.331 -0.274 -0.204 -0.256 -0.253
FedPxN
Method Metric 2 Clients 3 Clients 4 Clients 5 Clients Overall
Cosine Spearman 0.476 0.466 0.446 0.262 0.429
Kendall 0.339 0.324 0.307 0.182 0.298
Euclidean Spearman -0.509 -0.459 -0.435 -0.236 -0.425
Kendall -0.360 -0.317 -0.293 -0.168 -0.291
KL Spearman -0.507 -0.482 -0.395 -0.220 -0.416
Kendall -0.364 -0.346 -0.271 -0.154 -0.290
Appendix E Standard Deviation Results for Selecting Participating Subjects using Averaged Patient Embedding Similarity
Table 10: Standard Deviations for Table 4. The values represent standard deviations computed across three random seeds, each with distinct data splits and model initializations.
2 Average 0.004 0.004 0.004 0.006 0.004
Cosine 0.006 0.005 0.004 0.009 0.004
Euclidean 0.006 0.005 0.004 0.009 0.004
KL 0.006 0.005 0.004 0.009 0.004
Best 0.003 0.005 0.004 0.009 0.004
3 Average 0.004 0.004 0.003 0.005 0.005
Cosine 0.002 0.005 0.002 0.005 0.005
Euclidean 0.002 0.005 0.002 0.005 0.005
KL 0.002 0.005 0.002 0.005 0.005
Best 0.003 0.005 0.002 0.005 0.006
4 Average 0.003 0.005 0.002 0.006 0.006
Cosine 0.004 0.005 0.002 0.006 0.007
Euclidean 0.004 0.005 0.002 0.006 0.007
KL 0.004 0.005 0.002 0.006 0.007
Best 0.004 0.006 0.003 0.006 0.006
5 - 0.006 0.005 0.002 0.006 0.007
Appendix F Selecting Participating Subjects using Averaged Patient Embedding Similarity vs. Patient-level Embedding Similarity
Tables 11 and 12 present the host performance for each federated learning algorithm (i.e., FedAvg, FedProx, FedBN, FedPxN) when selecting participating clients using our averaged patient embedding-based similarity method. The results show that this method achieves performance equal to or better than using all five clients across all similarity metrics (i.e., cosine similarity, Euclidean distance, and KL divergence) in 9, 14, 8, and 12 out of 15 cases for FedAvg, FedProx, FedBN, and FedPxN, respectively. These values are computed by averaging the number of cases (across the three similarity metrics) where the performance is equal to or better than that of full five-client participation.
As discussed in Section 2.2.1, a potential concern with averaging patient embeddings across clients is the risk of information loss, which could lead to suboptimal client selection. To address this concern, we further evaluate the performance using patient-level embeddings similarities for selecting participating clients. Tables 13 and 14 present the host performance for each federated learning algorithm (i.e., FedAvg, FedProx, FedBN, FedPxN) when selecting participating clients using patient-level embedding similarities. Specifically, for Cosine Similarity, we compute host-subject patient-level similarities and then average the patient-level similarities to obtain the final similarity score between the host and subject. For Euclidean Distance, we similarly compute host-subject patient-level euclidean distances and then average the patient-level distances to obtain the final distance between the host and subject. For Precision, we follow the method in [27], computing precision based on individual data points within the host and subject embeddings. The results show that this method achieves performance equal to or better than using all five clients across all similarity metrics (i.e., cosine similarity, Euclidean distance, and KL divergence) in 7.67, 12.67, 7.33, and 9.33 out of 15 cases for FedAvg, FedProx, FedBN, and FedPxN, respectively. These values are computed by averaging the number of cases (across the three similarity metrics) where the performance is equal to or better than that of full five-client participation.
Overall, the results demonstrate that the averaged embedding approach is more robust and consistently achieves performance equal to or better than full participation, compared to the patient-level method. This highlights the efficacy and reliability of using averaged embeddings for selecting suitable clients for the host in federated learning.
Table 11: Host performance (average macro AUROC across 12 tasks) using the FedAvg and FedProx algorithms, when selecting participating clients based on averaged patient embeddings. The performance is compared to (1) using all 5 clients, (2) Average, and (3) Best. Average and Best refer to the average and best host performance, respectively, across all subject combinations for a fixed number of participating clients. Underlined values indicate performance equal to or better than when using all 5 clients, bold indicates performance equal to Best performance, asterisk indicates performance exceeding Average performance.
FedAvg
2 Average 0.806 0.791 0.814 0.779 0.774
Cosine 0.812* 0.797* 0.815* 0.783* 0.777*
Euclidean 0.812* 0.797* 0.815* 0.783* 0.777*
KL 0.812* 0.797* 0.815* 0.783* 0.777*
Best 0.812 0.797 0.815 0.783 0.777
3 Average 0.808 0.794 0.813 0.780 0.776
Cosine 0.813* 0.797* 0.814* 0.781* 0.777*
Euclidean 0.813* 0.797* 0.814* 0.781* 0.778*
KL 0.813* 0.797* 0.814* 0.781* 0.777*
Best 0.813 0.797 0.816 0.786 0.778
4 Average 0.810 0.793 0.813 0.782 0.776
Cosine 0.812* 0.793 0.811 0.786* 0.778*
Euclidean 0.812* 0.793 0.811 0.786* 0.778*
KL 0.812* 0.793 0.811 0.786* 0.778*
Best 0.812 0.797 0.816 0.786 0.778
5 - 0.812 0.796 0.812 0.782 0.780
FedProx
2 Average 0.804 0.792 0.812 0.779 0.772
Cosine 0.808* 0.800* 0.813* 0.784* 0.777*
Euclidean 0.808* 0.800* 0.813* 0.784* 0.777*
KL 0.808* 0.800* 0.813* 0.784* 0.777*
Best 0.811 0.800 0.813 0.784 0.777
3 Average 0.808 0.793 0.812 0.780 0.775
Cosine 0.814* 0.800* 0.814* 0.780 0.776*
Euclidean 0.814* 0.800* 0.814* 0.780 0.780*
KL 0.814* 0.800* 0.814* 0.780 0.776*
Best 0.814 0.800 0.816 0.784 0.780
4 Average 0.809 0.793 0.812 0.782 0.776
Cosine 0.813* 0.794 0.813 0.785* 0.778*
Euclidean 0.813* 0.794 0.813 0.785* 0.778*
KL 0.813* 0.794 0.813 0.785* 0.778*
Best 0.813 0.796 0.815 0.785 0.778
5 - 0.810 0.792 0.811 0.780 0.773
Table 12: Host performance (average macro AUROC across 12 tasks) using the FedBN and FedPxN algorithms, when selecting participating clients based on averaged patient embeddings. The performance is compared to (1) using all 5 clients, (2) Average, and (3) Best. Average and Best refer to the average and best host performance, respectively, across all subject combinations for a fixed number of participating clients. Underlined values indicate performance equal to or better than when using all 5 clients, bold indicates performance equal to Best performance, asterisk indicates performance exceeding Average performance.
FedBN
2 Average 0.808 0.793 0.814 0.779 0.775
Cosine 0.811* 0.798* 0.816* 0.785* 0.779*
Euclidean 0.811* 0.798* 0.816* 0.785* 0.779*
KL 0.811* 0.798* 0.816* 0.785* 0.779*
Best 0.811 0.798 0.816 0.785 0.779
3 Average 0.808 0.795 0.813 0.782 0.777
Cosine 0.812* 0.800* 0.816* 0.784* 0.777
Euclidean 0.812* 0.800* 0.816* 0.784* 0.777
KL 0.812* 0.800* 0.816* 0.784* 0.777
Best 0.812 0.800 0.816 0.787 0.779
4 Average 0.812 0.796 0.814 0.783 0.778
Cosine 0.815* 0.799* 0.813 0.787* 0.779*
Euclidean 0.815* 0.799* 0.813 0.787* 0.779*
KL 0.815* 0.799* 0.813 0.787* 0.779*
Best 0.815 0.799 0.816 0.787 0.779
5 Average 0.815 0.798 0.815 0.785 0.780
FedPxN
2 Average 0.807 0.794 0.812 0.779 0.774
Cosine 0.808* 0.800* 0.814* 0.785* 0.777*
Euclidean 0.808* 0.800* 0.814* 0.785* 0.777*
KL 0.808* 0.800* 0.814* 0.785* 0.777*
Best 0.812 0.800 0.815 0.785 0.777
3 Average 0.808 0.795 0.813 0.781 0.775
Cosine 0.812* 0.800* 0.813 0.786* 0.778*
Euclidean 0.812* 0.800* 0.813 0.786* 0.778*
KL 0.812* 0.800* 0.813 0.786* 0.778*
Best 0.813 0.800 0.815 0.786 0.779
4 Average 0.812 0.796 0.812 0.784 0.777
Cosine 0.815* 0.798* 0.812 0.788* 0.778*
Euclidean 0.815* 0.798* 0.812 0.788* 0.778*
KL 0.815* 0.798* 0.812 0.788* 0.778*
Best 0.815 0.799 0.816 0.788 0.778
5 Average 0.813 0.798 0.813 0.783 0.777
Table 13: Host performance (average macro AUROC across 12 tasks) using the FedAvg and FedProx algorithms, when selecting participating clients based on patient-level embeddings. The performance is compared to (1) using all 5 clients, (2) Average, and (3) Best. Average and Best refer to the average and best host performance, respectively, across all subject combinations for a fixed number of participating clients. Underlined values indicate performance equal to or better than when using all 5 clients, bold indicates performance equal to Best performance, asterisk indicates performance exceeding Average performance.
FedAvg
2 Average 0.806 0.791 0.814 0.779 0.774
Cosine 0.809 0.791 0.815* 0.774 0.774
Euclidean 0.809 0.791 0.815* 0.774 0.774
Precision 0.812* 0.797* 0.815* 0.783* 0.777*
Best 0.812 0.797 0.815 0.783 0.777
3 Average 0.808 0.794 0.813 0.780 0.776
Cosine 0.813* 0.797* 0.816* 0.781* 0.776
Euclidean 0.813* 0.797* 0.814* 0.776 0.776
Precision 0.813* 0.797* 0.814* 0.781* 0.777*
Best 0.813 0.797 0.816 0.786 0.778
4 Average 0.810 0.793 0.813 0.782 0.776
Cosine 0.812* 0.793 0.816* 0.786* 0.778*
Euclidean 0.812* 0.793 0.811 0.786* 0.772
Precision 0.812* 0.793 0.815* 0.786* 0.778*
Best 0.812 0.797 0.816 0.786 0.778
5 - 0.812 0.796 0.812 0.782 0.780
FedProx
2 Average 0.804 0.792 0.812 0.779 0.772
Cosine 0.811* 0.795* 0.813* 0.777 0.772
Euclidean 0.811* 0.795* 0.813* 0.777 0.772
Precision 0.808 0.800* 0.813* 0.784* 0.777*
Best 0.811 0.800 0.813 0.784 0.777
3 Average 0.808 0.793 0.812 0.780 0.775
Cosine 0.814* 0.800* 0.815* 0.780 0.776*
Euclidean 0.814* 0.800* 0.814* 0.777 0.776*
Precision 0.814* 0.800* 0.814* 0.780 0.776*
Best 0.814 0.800 0.816 0.784 0.780
4 Average 0.809 0.793 0.812 0.782 0.776
Cosine 0.813* 0.794* 0.815* 0.785* 0.778*
Euclidean 0.813* 0.794* 0.813 0.785* 0.776
Precision 0.813* 0.794* 0.810 0.785* 0.778*
Best 0.813 0.796 0.815 0.785 0.778
5 - 0.810 0.792 0.811 0.780 0.773
Table 14: Host performance (average macro AUROC across 12 tasks) using the FedBN and FedPxN algorithms, when selecting participating clients based on patient-level embeddings. The performance is compared to (1) using all 5 clients, (2) Average, and (3) Best. Average and Best refer to the average and best host performance, respectively, across all subject combinations for a fixed number of participating clients. Underlined values indicate performance equal to or better than when using all 5 clients, bold indicates performance equal to Best performance, asterisk indicates performance exceeding Average performance.
FedBN
2 Average 0.808 0.793 0.814 0.779 0.775
Cosine 0.811* 0.794* 0.816* 0.774 0.774
Euclidean 0.811* 0.794* 0.816* 0.774 0.774
Precision 0.811* 0.798* 0.816* 0.785* 0.779*
Best 0.811 0.798 0.816 0.785 0.779
3 Average 0.808 0.795 0.813 0.782 0.777
Cosine 0.812* 0.800* 0.815* 0.784* 0.774
Euclidean 0.812* 0.800* 0.816* 0.779 0.774
Precision 0.812* 0.800* 0.816* 0.784* 0.777
Best 0.812 0.800 0.816 0.787 0.779
4 Average 0.812 0.796 0.814 0.783 0.778
Cosine 0.815* 0.799* 0.816* 0.787* 0.779*
Euclidean 0.815* 0.799* 0.813 0.787* 0.778
Precision 0.815* 0.799* 0.815* 0.787* 0.779*
Best 0.815 0.799 0.816 0.787 0.779
5 - 0.815 0.798 0.815 0.785 0.780
FedPxN
2 Average 0.807 0.794 0.812 0.779 0.774
Cosine 0.812* 0.794* 0.814* 0.777* 0.775*
Euclidean 0.812* 0.794* 0.814* 0.777* 0.775*
Precision 0.808* 0.800* 0.814* 0.785* 0.777*
Best 0.812 0.800 0.815 0.785 0.777
3 Average 0.808 0.795 0.813 0.781 0.775
Cosine 0.812* 0.800* 0.815* 0.786* 0.775
Euclidean 0.812* 0.800* 0.813 0.776 0.775
Precision 0.812* 0.800* 0.813 0.786* 0.778*
Best 0.813 0.800 0.815 0.786 0.779
4 Average 0.812 0.796 0.812 0.784 0.777
Cosine 0.815* 0.798* 0.816* 0.788* 0.778*
Euclidean 0.815* 0.798* 0.812 0.788* 0.777
Precision 0.815* 0.798* 0.812 0.788* 0.778*
Best 0.815 0.799 0.816 0.788 0.778
5 - 0.813 0.798 0.813 0.783 0.777

### v5

@@ 4.2 Relationship between Averaged Patient Embedding Similarity and Host Performance
Table 5: Correlation between host-subject similarity and the change in host performance when a subject participates in federated learning compared to when the subject does not. The correlations are averaged across that of FedAvg, FedProx, FedBN, and FedPxN. Similarity is measured using averaged patient embeddings (or averaged softmax for KL divergence). See Supplementary Table S5 for details.
@@
Table 5 presents the observed correlations between (1) similarity between the host and a subject, measured using averaged patient embeddings (or averaged softmax values for KL Divergence) and (2) the difference in host performance with and without the subject participating in federated learning (if multiple subjects participate, other subjects consistently participate in both scenarios to isolate the effect of the target subject). The results show consistent correlations across all three similarity methods and the different numbers of participating clients. This trend holds across each of the four federated learning algorithms (see Supplementary Table S5). Note that Cosine Similarity exhibits positive correlations (as higher values indicate greater similarity), while Euclidean Distance and KL Divergence exhibit negative correlations (as higher values indicate lower similarity). These correlations indicate that averaged patient embedding similarity can serve as a reliable criterion for selecting which subjects to exclude from federated learning, particularly by eliminating those with lower similarity scores (or larger distances).
@@ 4.3 Selecting Participating Subjects using Averaged Patient Embedding Similarity
Table 6: Host performance (average macro AUROC across 12 tasks) when selecting participating subjects using averaged patient embeddings (Cosine Sim., Euclidean Dist., and KL Div.), averaged across FedAvg, FedProx, FedBN, FedPxN. The performance is compared to (1) using all 5 clients, (2) Average, and (3) Best. Average and Best refer to the average and best host performance, respectively, across all subject combinations for a fixed number of participating clients. Underlined values indicate performance equal to or better than when using all 5 clients, bold indicates performance equal to Best performance, asterisk indicates performance exceeding Average performance. Standard deviations are listed in Supplementary Table S6.
@@
Detailed performance results for each individual federated learning algorithm are presented in Supplementary Tables S7 and S8. These results show that our selection method consistently achieves performance equal to or better than using all 5 clients and consistently matches the Best and exceeds the Average performance in most cases across each of the four algorithms. Overall, these results highlight the effectiveness of our averaged patient embedding-based method in identifying suitable subjects for the host, thereby enabling effective reduction in the number of participating clients without sacrificing host model performance.
@@ Discussion
In this work, we address two practical challenges that arise when building federated learning models tailored to a single healthcare institution: (1) heterogeneity across EHR systems and (2) budget constraints associated with building the federated learning model. As highlighted in prior work [2], differences in EHR schemas and medical coding systems pose a fundamental barrier to collaborative modeling across healthcare institutions. At the same time, recent studies emphasize that cost considerations are critical for real-world adoption of modeling in healthcare, particularly under constrained institutional budgets [10, 11]. Our proposed EHRFL framework jointly addresses these challenges by enabling compatible text-based EHR modeling while introducing a cost-aware participant selection strategy in federated learning.
Assumptions
As is common in healthcare institutions, our framework is designed for a cross-silo federated learning setting, where participating institutions are assumed to have relatively stable network connectivity. Under this assumption, the communication of model parameters and averaged embeddings does not constitute a major bottleneck, in contrast to cross-device federated learning scenarios characterized by unstable or bandwidth-limited connections.
Criteria for Selecting Suitable Participating Clients
One of the core components of our framework is the use of averaged patient embedding similarity as a criterion for selecting suitable federated learning participants. Our experimental results show that using our averaged patient embedding similarity method enables identifying and excluding subjects that contribute marginally or negatively, effectively reducing federated learning costs without compromising performance. The use of embedding-based dataset representations aligns with prior research on dataset similarity and representation learning [36, 37, 38], while extending these ideas to a privacy-preserving, federated healthcare setting. Although embedding similarity is effective in our experiments, alternative criteria—such as label distributions, feature statistics, or performance-based metrics—could also inform participant selection [39, 40, 41]. A systematic comparison of different selection criteria may further improve robustness, which we leave for future work as this is the first work for participating client selection.
Scalability of Selecting Suitable Participating Clients
Our participant selection strategy requires training the embedding model only once, which is performed by the host. The trained model weights are then shared with candidate subjects, who use the model solely for embedding extraction and do not perform any model training. As a result, the per-subject overhead is limited to lightweight inference for embedding extraction and the communication of a single averaged embedding. Even when the number of candidate institutions is large, the dominant computational cost—model training—remains constant, while communication and inference costs scale linearly and remain modest.
Limitations and Future Works
A limitation of our work is that our experiments is restricted to ICU datasets, due to resource constraints. Although we incorporate multiple heterogeneous ICU datasets spanning different EHR systems, further validation on other clinical settings—such as outpatient, emergency department data—can be important to assess the generalizability of our approach. In addition, our framework does not explicitly incorporate data quality checks or data cleaning procedures during the data preprocessing stage. Integrating data quality control and data cleaning strategies, as studied in prior research [42, 43, 44], may help remove erroneous data and further improve the robustness and reliability of the text representations. We leave such applications to future work.
Broader Applicability beyond Healthcare
Beyond healthcare, the principles underlying EHRFL may generalize to other domains in which a single institution aims to build a predictive model for their needs with other institutions under schema heterogeneity and cost constraints. For example, in retail and e-commerce forecasting, organizations often operate heterogeneous point-of-sale and inventory systems [45], and participant selection based on consumer behavioral similarity could reduce collaboration costs without exposing sensitive transaction data. Similarly, in manufacturing and supply-chain predictive maintenance, factories often use different Industrial IoT platforms and sensor systems. Sensor platforms and maintenance logs could be linearized into text format regardless of the underlying data schema, and manufacturers could select federated learning participants with similar operating conditions. These extensions suggest that the proposed framework may have broader applicability beyond clinical prediction tasks.
Conclusion
@@
In this paper, we introduce EHRFL, a federated learning framework for developing a client-specific model using EHRs. By incorporating text-based EHR modeling, EHRFL enables federated learning across clients with heterogeneous EHR systems. Furthermore, our framework optimizes the selection of participating clients using averaged patient embeddings, reducing the number of participants without compromising model performance.
This work can benefit multiple stakeholders in the healthcare ecosystem. Healthcare institutions can leverage EHRFL to build predictive models tailored to their internal needs while controlling federated learning costs and preserving patient privacy. Clinical data departments and can adopt the text-based EHR federated learning approach to enable collaborative modeling across incompatible EHR systems with minimal preprocessing overhead. From a research perspective, EHRFL provides a practical framework for studying federated learning under realistic healthcare constraints, including heterogeneity, privacy, and budget limitations. We release open-source code and experimental details to support reproducibility and adoption.
As next steps, we plan to extend EHRFL in several directions. First, we aim to evaluate the framework on additional clinical settings beyond the ICU, such as outpatient and emergency department data, to further assess generalizability. Second, we plan to incorporate data quality control mechanisms prior to text-based linearization to improve robustness in real-world deployments. Third, we will explore alternative client selection criteria to compare and enhance the selection robustness. These extensions can be addressed incrementally in future work, with a focus on improving generalizability, robustness, and evaluation breadth. We believe these steps provide a sustainable path toward deploying cost-effective federated learning systems in practical healthcare environments.